# Pullback of arithmetic theta series and its modularity for unitary Shimura curves

### v1

@@ 3.1. Extensions of the line bundles to the boundary
Let $\mathcal{Z}_{p}^{*}$ be the Zariski closure of $\mathcal{Z}_{p}$ in $\mathcal{S}_{p}^{*}$, then $\mathcal{Z}_{p}^{*}=\mathcal{S}_{p}^{*}-\hbox{Exc}_{p}$ is a Cartier divisor of $\mathcal{S}_{p}^{*}$. Since $\mathcal{S}_{p}^{*}$ is the principal divisor defined by $p$, we have the following equation in $\mathrm{CH}^{1}(\mathcal{S}^{*})$.
@@
By [${}^{+}$20, Theorem 3.7.1(6)], the line bundle $\boldsymbol{\omega}$ admits a canonical extension to the compactification $\mathcal{S}^{*}$ which is still denoted by $\boldsymbol{\omega}$ such that
(3.8) $\boldsymbol{\omega}^{-1}|_{z}=\operatorname{Hom}(\mathrm{Fil}(\mathbf{A}_{0}),\operatorname{Lie}(\mathbf{G}_{z})/\mathcal{F}_{\mathbf{G}_{z}}).$
@@
Now let $z$ be any geometric close point of $\mathcal{S}\setminus\mathrm{Exc}$. Since $L_{\mathbf{A}}$ has rank $1$, if the inclusion $\bar{\epsilon}\operatorname{Lie}{\mathbf{A}_{z}}\subset L_{\mathbf{A}_{z}}$ is strict, we must have $\bar{\epsilon}\operatorname{Lie}{\mathbf{A}_{z}}=0$. In other words $\bar{\epsilon}\operatorname{Lie}\mathbf{A}_{z}=\{0\}$, which forces $z$ to be a point in $\mathrm{Exc}$ by the definition of $\mathrm{Exc}$. This is a contradiction. So we know that the lemma is true on the level of geometric close points. Since $\bar{\epsilon}:\operatorname{Lie}\mathbf{A}\rightarrow\operatorname{Lie}\mathbf{A}$ is a section of the coherent sheaf $\operatorname{End}(\operatorname{Lie}\mathbf{A})$, the lemma is true by Nakayama’s lemma. ∎
@@
We first claim that $\bar{\epsilon}\operatorname{Lie}{\mathbf{G}}$ is a line bundle over $\mathcal{S}^{*}\setminus\mathrm{Exc}$. By Nakayama’s lemma, it suffices to check all geometric close points. For points in $\mathcal{S}\setminus\mathrm{Exc}$, the claim follows from Lemma 3.6. For a geometric point $z$ on $\mathcal{B}_{\Phi}$, we have the exact sequence of Lie algebras
@@
The argument in Step 1 of the proof of Proposition 3.5 extends to the boundary. Hence none of the boundary component belongs to $(s)$. Since both sides of the equation are Cartier divisors, the corollary follows from Proposition 3.5. ∎
@@ 3.2. Constant term of the generating series
By Theorem 3.2, the definition of $\mathcal{Z}^{*}(0)$ (the constant term in the generating series of special divisors) agrees with that of [${}^{+}$20, Equation (1.1.4)]. We will see in the next section that as $\Omega$ controls the deformation of special divisors, this definition of the constant term is also compatible with pullbacks, see §5.
4. Deformation of special divisors
It is discovered by [How19] that the line bundle $\Omega$ controls the deformation theory of special divisors on the open Shimura variety $\mathcal{S}$. The main goal of this section is to extend this observation to $\mathcal{S}^{*}$. First we review some deformation theory that we will need following [How19, §4]. Throughout this section we will repeatedly encounter the following setting. Let $S$ be a stack and $Z$ be another stack with a finite unramified morphism $Z\rightarrow S$. We assume there is an étale morphism of stacks $U\rightarrow S$ such that $Z|_{U}\rightarrow S|_{U}$ is a closed immersion of schemes on every connected component of its domain. Let $\mathcal{Z}$ be such a connected component and $I_{\mathcal{Z}}$ be its ideal sheaf in $S|_{U}$. Then $I_{\mathcal{Z}}^{2}$ defines a larger closed formal subscheme
@@ 4. Deformation of special divisors
$H_{1}^{\mathrm{dR}}(\mathbf{A}_{0}|_{\tilde{\mathcal{Z}}})\xrightarrow{\tilde{x}}H_{1}^{\mathrm{dR}}(\mathbf{A}|_{\tilde{\mathcal{Z}}}),$
@@
(4.1) $\mathrm{Fil}(\mathbf{A}_{0}|_{\tilde{\mathcal{Z}}})\xrightarrow{\tilde{x}}\operatorname{Lie}(\mathbf{A}|_{\tilde{\mathcal{Z}}}).$
@@
The Kudla-Rapoport divisor $\mathcal{Z}$ is the largest closed formal subscheme of $\tilde{\mathcal{Z}}$ over which (4.2) is trivial.
@@
We will extend Proposition 4.1 to the toroidal compactification $\mathcal{S}^{*}$, see Theorem 4.4.
4.1. Special divisors on boundary charts
We would like to study special divisors on the boundary chart $\mathcal{C}_{\Phi}^{*}$. As in §2.4, assume $\Phi=(\mathfrak{n},L_{\Phi})$ and $\mathfrak{m}$ be the hermitian dual of $\mathfrak{n}$. First let us construct a semi-abelian scheme $G^{\sharp}$ over $\mathcal{B}_{\Phi}$ and a $1$-motive $M$ over $\mathcal{C}_{\Phi}$ as follows.
By the definition of $\mathcal{B}_{\Phi}$, we get a tautological homomorphism of fppf sheaves $c^{\vee}\in\underline{\operatorname{Hom}}_{\mathcal{O}_{\boldsymbol{k}}}(\mathfrak{n},B)$ over $\mathcal{B}_{\Phi}$. Since
(see for example [Lan13, Proposition 3.1.5.1]), $c^{\vee}$ determines a semi-abelian scheme $(G^{\sharp})^{\vee}$ over $\mathcal{B}_{\Phi}$, such that there is an exact sequence of fppf sheaves of $\mathcal{O}_{\boldsymbol{k}}$-modules
(4.3) $0\rightarrow\mathfrak{m}\otimes_{\mathbb{Z}}\mathbb{G}_{m}\rightarrow(G^{\sharp})^{\vee}\xrightarrow{\pi^{\prime}}B^{\vee}\rightarrow 0$
Similarly define $c=\psi\circ c^{\vee}$ where $\psi:B\rightarrow B^{\vee}$ is the principal polarization of $B$, then $c$ determines a semi-abelian scheme $G^{\sharp}$ over $\mathcal{B}_{\Phi}$, such that there is an exact sequence of fppf sheaves of $\mathcal{O}_{\boldsymbol{k}}$-modules
[…]
We remark here that $G^{\sharp}$ agree with the pullback of $\mathbf{G}|_{\mathcal{S}^{*}(\Phi)}$ under the morphism $\pi_{\Phi}$ in Theorem 2.5 (3). By [Lan13, Lemma 3.4.2], the condition $c=\psi\circ c^{\vee}$ guarantees that there is an $\mathcal{O}_{\boldsymbol{k}}$-linear isomorphism $\lambda^{\sharp}:G^{\sharp}\rightarrow(G^{\sharp})^{\vee}$. Since $\mathcal{C}_{\Phi}$ and $\mathcal{C}_{\Phi}^{*}$ are $\mathcal{B}_{\Phi}$-stacks, we can pullback $(G^{\sharp})^{\vee}$ and $G^{\sharp}$ to semi-abelian schemes on $\mathcal{C}_{\Phi}$ and $\mathcal{C}_{\Phi}^{*}$ (denoted by the same notations).
The definition of $\mathcal{C}_{\Phi}$ and $\mathcal{C}_{\Phi}^{*}$ gives us in addition a positive, symmetric, and $\mathcal{O}_{\boldsymbol{k}}$-linear isomorphism
$\tau:1_{(\underline{\mathfrak{n}}\times\underline{\mathfrak{n}})|_{\mathcal{C}_{\Phi}}}\rightarrow(c^{\vee}\times c)^{*}(\mathcal{P}^{-1})|_{(\underline{\mathfrak{n}}\times\underline{\mathfrak{n}})|_{\mathcal{C}_{\Phi}}}$
[…]
$\pi\circ u=c^{\vee},\ \pi^{\prime}\circ v=c.$
By the proof of [${}^{+}$20, Proposition 3.3.3], the morphism $M\rightarrow M^{\vee}$ induced by the identity map of $\mathfrak{n}$ and $\lambda^{\sharp}$ is a principal polarization of $M$ in the sense of [Del74, §10.2.11], which is compatible with the given polarization $\psi:B\rightarrow B^{\vee}$ and with the isomorphism $\mathfrak{m}\cong\mathfrak{n}^{\vee}$.
Suppose we have a morphism of group schemes $x:\mathbf{A}_{0}\rightarrow G^{\sharp}$ over $\mathcal{B}_{\Phi}$. Its dual morphism $x^{\vee}:(G^{\sharp})^{\vee}\rightarrow\mathbf{A}_{0}^{\vee}$ is defined as follows. By the definition of the dual $1$-motive $M^{\vee}$, we have
@@ 4.1. Special divisors on boundary charts
[…]
If we view $B^{\vee}$ as $\underline{\mathrm{Ext}}^{1}(B,\mathbb{G}_{m})$, we immediately see that $x^{\vee}=(\pi\circ x)^{\vee}\circ\pi^{\prime}$ where $\pi$ and $\pi^{\prime}$ are defined in (4.4) and (4.3). Now we are ready to define special divisors on $\mathcal{B}_{\Phi}$, $\mathcal{C}_{\Phi}$ and $\mathcal{C}_{\Phi}^{*}$. The following definition is due to [${}^{+}$20, Remark 3.6.3].
@@
For $m\in\mathbb{Z}_{>0}$, define $\mathcal{Z}_{\mathcal{B}_{\Phi}}(m)$ to be the subscheme of $\mathcal{B}_{\Phi}$ such that given a morphism of $\mathcal{O}_{\boldsymbol{k}}$-scheme $S\rightarrow\mathcal{C}_{\Phi}$,
$\mathcal{Z}_{\mathcal{B}_{\Phi}}(m)(S)=\{x\in\operatorname{Hom}_{\mathcal{O}_{\boldsymbol{k}}}(\mathbf{A}_{0,S}\ ,G^{\sharp})\mid\langle x,x\rangle=m\}.$
Here $\langle$,$\rangle$ , ⟩ is defined as in (2.9):
(4.5) $\left\langle x_{1},x_{2}\right\rangle=\iota_{0}^{-1}(\lambda_{0}^{-1}\circ x_{2}^{\vee}\circ\lambda^{\sharp}\circ x_{1})\in\mathcal{O}_{\boldsymbol{k}}$
@@
By [${}^{+}$20, Proposition 3.6.2], $\mathcal{Z}_{\mathcal{B}_{\Phi}}(m)$ is an effective Cartier divisor of $\mathcal{B}_{\Phi}$ that is flat over $\mathcal{A}_{\Phi}$ and $\mathcal{O}_{\boldsymbol{k}}$. As $\mathcal{C}_{\Phi}$ and $\mathcal{C}_{\Phi}^{*}$ are smooth over $\mathcal{B}_{\Phi}$, $\mathcal{Z}_{\Phi}(m)$ and $\mathcal{Z}_{\Phi}^{*}(m)$ are smooth over $\mathcal{Z}_{\mathcal{B}_{\Phi}}(m)$, and are effective Cartier divisors of $\mathcal{C}_{\Phi}$ and $\mathcal{C}_{\Phi}^{*}$ respectively, flat over $\mathcal{A}_{\Phi}$ and $\mathcal{O}_{\boldsymbol{k}}$.
4.2. Deformation of special divisors on boundary charts
Now consider $G^{\sharp}$ as the $1$-motive $[0\rightarrow G^{\sharp}]$ over $\mathcal{B}_{\Phi}$. Let $H_{1}^{\mathrm{dR}}(G^{\sharp})$ be the De Rham realization of $G^{\sharp}$ (see [Del74, §10.1.7]) defined by the Lie algebra of its universal vector extension, which is a rank $22n-$ 2 vector bundle over $\mathcal{B}_{\Phi}$, locally free of rank $n-1$ over $\mathcal{O}_{\boldsymbol{k}}\otimes_{\mathbb{Z}}\mathcal{O}_{\mathcal{B}_{\Phi}}$. Then $\mathrm{Fil}(G^{\sharp})=F^{0}H_{1}^{\mathrm{dR}}(G^{\sharp})$ is a locally free direct summand of $H_{1}^{\mathrm{dR}}(G^{\sharp})$ of $\mathcal{O}_{\mathcal{B}_{\Phi}}$-rank $n-2$. We have the following exact sequence of Lia algebras
(4.7) $0\rightarrow\mathfrak{m}\otimes_{\mathbb{Z}}\mathcal{O}_{\mathcal{B}_{\Phi}}\rightarrow\operatorname{Lie}G^{\sharp}\rightarrow\operatorname{Lie}B\rightarrow 0$

### v2

@@ 3.1. Extensions of the line bundles to the boundary
Assume $p$ is a prime divisor of $D$. Let $\mathcal{Z}_{p}^{*}$ be the Zariski closure of $\mathcal{Z}_{p}$ in $\mathcal{S}_{p}^{*}$, then $\mathcal{Z}_{p}^{*}+\hbox{Exc}_{p}=\mathcal{S}_{p}^{*}$ as Cartier divisors on $\mathcal{S}^{*}$. Since $\mathcal{S}_{p}^{*}$ is the Cartier divisor defined by an ideal $\mathfrak{p}$ such that $\mathfrak{p}^{2}=(p)$, we have the following equation in $\mathrm{CH}_{\mathbb{Q}}^{1}(\mathcal{S}^{*})$.
@@
By [BHK+20, Theorem 3.7.1(6)], the line bundle $\boldsymbol{\omega}$ admits a canonical extension to the compactification $\mathcal{S}^{*}$ which is still denoted by $\boldsymbol{\omega}$ such that
(3.10) $\boldsymbol{\omega}^{-1}=\operatorname{Hom}(\mathrm{Fil}(\mathbf{A}_{0}),\operatorname{Lie}(\mathbf{G})/\mathcal{F}_{\mathbf{G}}).$
@@
Now let $z$ be any geometric closed point of $\mathcal{S}\setminus\mathrm{Exc}$. Since $L_{\mathbf{A}}$ has rank $1$, if the inclusion $\bar{\epsilon}\operatorname{Lie}{\mathbf{A}_{z}}\subset L_{\mathbf{A}_{z}}$ is strict, we must have $\bar{\epsilon}\operatorname{Lie}{\mathbf{A}_{z}}=0$. In other words $\bar{\epsilon}\operatorname{Lie}\mathbf{A}_{z}=\{0\}$, which forces $z$ to be a point in $\mathrm{Exc}$ by the definition of $\mathrm{Exc}$. This is a contradiction. So we know that the lemma is true on the level of geometric closed points. Since $\bar{\epsilon}:\operatorname{Lie}\mathbf{A}\rightarrow\operatorname{Lie}\mathbf{A}$ is a section of the coherent sheaf $\operatorname{End}(\operatorname{Lie}\mathbf{A})$, the lemma is true by Nakayama’s lemma. ∎
@@
We first claim that $\bar{\epsilon}\operatorname{Lie}{\mathbf{G}}$ is a line bundle over $\mathcal{S}^{*}\setminus\mathrm{Exc}$. By Nakayama’s lemma, it suffices to check all geometric closed points. For points in $\mathcal{S}\setminus\mathrm{Exc}$, the claim follows from Lemma 3.6. For a geometric point $z$ on $\mathcal{B}_{\Phi}$, we have the exact sequence of Lie algebras
@@
The flatness of the boundary divisor $\partial\mathcal{S}^{*}$ implies that every irreducible component of $\partial\mathcal{S}^{*}$ has some closed point of characteristic prime to $D$, and Step 1 of the proof of Proposition 3.5 implies that such a point cannot lie in $(s)$. Hence none of the boundary component belongs to $(s)$. Since both sides of the equation are Cartier divisors, the corollary follows from Proposition 3.5. ∎
@@ 3.2. Constant term of the generating series
By Theorem 3.2, the definition of $\mathcal{Z}^{*}(0)$ (the constant term in the generating series of special divisors) agrees with that of [BHK+20, Equation (1.1.4)]. We will see in the next section that as $\Omega$ controls the deformation of special divisors, this definition of the constant term is also compatible with pullbacks, see §5.
4. Deformation of special divisors near the boundary
It is discovered by [How19] that the line bundle $\Omega$ controls the deformation theory of special divisors on the open Shimura variety $\mathcal{S}$, see Proposition 4.1 below. The main goal of this section is to extend this observation to $\mathcal{S}^{*}$ (Proposition 4.8). In order to do this we define $\mathcal{Z}^{*}(m)$ as a Deligne-Mumford stack (see Definition 4.5 and Proposition 4.7), which should be thought of as the “toroidal compactification” of the stack $\mathcal{Z}(m)$. By Proposition 4.9 and [BHK+20, Theorem 3.7.1(4)], our definition of $\mathcal{Z}^{*}(m)$ agrees with that of [BHK+20] which is the Zariski closure of $\mathcal{Z}(m)$.
First we review some deformation theory following [How19, §4]. Throughout this section we will repeatedly encounter the following setting. Let $S$ be a stack and $Z$ be another stack with a finite unramified morphism $Z\rightarrow S$. So there is an étale morphism $U\rightarrow S$ such that $U$ is a scheme and $Z|_{U}\rightarrow S|_{U}$ is a closed immersion of schemes on every connected component of its domain. Let $\mathcal{Z}$ be such a connected component and $I_{\mathcal{Z}}$ be its ideal sheaf in $S|_{U}$. Then $I_{\mathcal{Z}}^{2}$ defines a larger closed subscheme
@@ 4. Deformation of special divisors near the boundary
$H_{1}^{\mathrm{dR}}(\mathbf{A}_{0}|_{\mathcal{Z}})\xrightarrow{x}H_{1}^{\mathrm{dR}}(\mathbf{A}|_{\mathcal{Z}}),$
@@
[…]
@@
The Kudla-Rapoport divisor $\mathcal{Z}$ is the largest closed subscheme of $\tilde{\mathcal{Z}}$ over which (4.2) is trivial.
@@
4.1. Special divisors on formal boundary charts
We would like to study special divisors on the boundary chart $\mathcal{C}_{\Phi}^{*}$. Let $\Phi=(\mathfrak{n},L_{\Phi})$ and $\mathfrak{m}$ be as in §2.7. We are in the setting of §2.5 with
$(X,Z,U)=(\mathcal{C}_{\Phi}^{*},\mathcal{B}_{\Phi},\mathcal{C}_{\Phi}).$
Recall from §2.5 and §2.6 that there is tautological degeneration data $(A_{0},\iota_{0},\lambda_{0},B,\kappa,\psi,c,c^{\vee},\tau)$ of type $\Phi$ and signature (n1,1)(n-1,1) over $(\mathcal{C}_{\Phi}^{*},\mathcal{B}_{\Phi},\mathcal{C}_{\Phi})$, and using this we can construct semi-abelian schemes $G^{\sharp}$ and $(G^{\sharp})^{\vee}$ and $1$-motives $M$ over $M^{\vee}$ over $\mathcal{C}_{\Phi}^{*}$. In fact, $G^{\sharp}$ and $(G^{\sharp})^{\vee}$ are defined over $\mathcal{B}_{\Phi}$ and can be viewed as objects over $\mathcal{C}_{\Phi}^{*}$ through the canonical projection $\mathcal{C}_{\Phi}^{*}\rightarrow\mathcal{B}_{\Phi}$.
Suppose we have a morphism of group schemes $x:A_{0}\rightarrow G^{\sharp}$ over $\mathcal{B}_{\Phi}$. Its dual morphism $x^{\vee}:(G^{\sharp})^{\vee}\rightarrow A_{0}^{\vee}$ is defined as follows. By the definition of the dual $1$-motive $M^{\vee}$, we have
@@ 4.1. Special divisors on formal boundary charts
$\tilde{A}_{0}$$A_{0}$$\tilde{B}$${B.}{\square}$$\scriptstyle{\pi\circ x}$
If we view $B^{\vee}$ as $\underline{\mathrm{Ext}}^{1}(B,\mathbb{G}_{m})$, we immediately see that
(4.3) $x^{\vee}=(\pi\circ x)^{\vee}\circ\pi^{\prime}$
where $\pi$ and $\pi^{\prime}$ are defined in (2.14) and (2.13). Now we are ready to define special divisors on $\mathcal{B}_{\Phi}$, $\mathcal{C}_{\Phi}$ and $\mathcal{C}_{\Phi}^{*}$. For any morphism of $\mathcal{O}_{\boldsymbol{k}}$-schemes $S\rightarrow\mathcal{B}_{\Phi}$, we define a hermitian form on $\operatorname{Hom}_{\mathcal{O}_{\boldsymbol{k}}}(A_{0,S}\ ,G^{\sharp}_{S})$ in the same way as in (2.9):
(4.4) $\left(x_{1},x_{2}\right)=\iota_{0}^{-1}(\lambda_{0}^{-1}\circ x_{2}^{\vee}\circ\lambda^{\sharp}\circ x_{1})\in\mathcal{O}_{\boldsymbol{k}}$
Lemma 4.2.
For any $x_{1},x_{2}\in\operatorname{Hom}_{\mathcal{O}_{\boldsymbol{k}}}(A_{0,S}\ ,G^{\sharp}_{S})$, we have
$\left(x_{1},x_{2}\right)=\left(\pi\circ x_{1},\pi\circ x_{2}\right)_{\psi},$
where $\pi:G^{\sharp}\rightarrow B$ is defined in (2.14) and $(,)_{\psi}$ is the hermitian form on $\operatorname{Hom}_{\mathcal{O}_{\boldsymbol{k}}}(A_{0,S}\ ,B_{S})$ defined in the same way as (2.9) using the polarization $\psi:B\rightarrow B^{\vee}$. In particular $\left(,\right)$ is positive definite.
Proof.
We have
$\left(x_{1},x_{2}\right)=$ $\iota_{0}^{-1}(\lambda_{0}^{-1}\circ x_{2}^{\vee}\circ\lambda^{\sharp}\circ x_{1})$
$=$ $\iota_{0}^{-1}(\lambda_{0}^{-1}\circ(\pi\circ x_{2})^{\vee}\circ\pi^{\prime}\circ\lambda^{\sharp}\circ x_{1})$
$=$ $\iota_{0}^{-1}(\lambda_{0}^{-1}\circ(\pi\circ x_{2})^{\vee}\circ\psi\circ\pi\circ x_{1})$
$=$ $\left(\pi\circ x_{1},\pi\circ x_{2}\right)_{\psi}.$
The third equality above uses the compatibility of $\psi$ and $\lambda^{\sharp}$: $\pi^{\prime}\circ\lambda^{\sharp}=\psi\circ\pi$ by the definition of polarization of $1$-motives.
If $\pi\circ x_{1}=0$, then $x_{1}$ factors through $\mathfrak{m}\otimes\mathbb{G}_{m}$ which is affine. This forces $x_{1}$ to be 0. The positivity of $\left(,\right)$ now follows from that of $\left(,\right)_{\psi}$ by [KR14, Lemma 2.7]. ∎
The following definition is due to [BHK+20, Remark 3.6.3].
@@
For $m\in\mathbb{Z}_{>0}$, define $\mathcal{Z}_{\mathcal{B}_{\Phi}}(m)$ to be the stack assigning for any morphism of $\mathcal{O}_{\boldsymbol{k}}$-schemes $S\rightarrow\mathcal{B}_{\Phi}$, the set
$\mathcal{Z}_{\mathcal{B}_{\Phi}}(m)(S)=\{x\in\operatorname{Hom}_{\mathcal{O}_{\boldsymbol{k}}}(A_{0,S}\ ,G^{\sharp}_{S})\mid(x,x)=m\}.$
@@
By [BHK+20, Proposition 3.6.2], the natural morphism $\mathcal{Z}_{\mathcal{B}_{\Phi}}(m)\rightarrow\mathcal{B}_{\Phi}$ is finite and unramified as $\mathcal{Z}_{\mathcal{B}_{\Phi}}(m)$ is a finite disjoint union of stacks, each of which maps to $\mathcal{B}_{\Phi}$ as a Cartier divisor. We can view $\mathcal{Z}_{\mathcal{B}_{\Phi}}(m)$ (or rather its image under the morphism to $\mathcal{B}_{\Phi}$) as an effective Cartier divisor of $\mathcal{B}_{\Phi}$ that is flat over $\mathcal{A}_{\Phi}$ and $\mathrm{Spec}\,\mathcal{O}_{\boldsymbol{k}}$. As $\mathcal{C}_{\Phi}$ and $\mathcal{C}_{\Phi}^{*}$ are smooth over $\mathcal{B}_{\Phi}$, $\mathcal{Z}_{\Phi}(m)$ and $\mathcal{Z}_{\Phi}^{*}(m)$ are smooth over $\mathcal{Z}_{\mathcal{B}_{\Phi}}(m)$, and are effective Cartier divisors of $\mathcal{C}_{\Phi}$ and $\mathcal{C}_{\Phi}^{*}$ respectively, flat over $\mathcal{A}_{\Phi}$ and $\mathrm{Spec}\,\mathcal{O}_{\boldsymbol{k}}$.
4.2. Deformation of special divisors on formal boundary charts
Now consider $G^{\sharp}$ as the $1$-motive $0\rightarrow G^{\sharp}]$ over $\mathcal{B}_{\Phi}$. Let $H_{1}^{\mathrm{dR}}(G^{\sharp})$ be the de Rham realization of $G^{\sharp}$ (see [Del74, §10.1.7]) defined by the Lie algebra of its universal vector extension, which is a rank 2n22n-2 vector bundle over $\mathcal{B}_{\Phi}$, locally free of rank n1n-1 over $\mathcal{O}_{\boldsymbol{k}}\otimes_{\mathbb{Z}}\mathcal{O}_{\mathcal{B}_{\Phi}}$. Then $\mathrm{Fil}(G^{\sharp})=F^{0}H_{1}^{\mathrm{dR}}(G^{\sharp})$ is a locally free direct summand of $H_{1}^{\mathrm{dR}}(G^{\sharp})$ of $\mathcal{O}_{\mathcal{B}_{\Phi}}$-rank n2n-2. We have the following exact sequence of Lie algebras
(4.6) $0\rightarrow\mathfrak{m}\otimes_{\mathbb{Z}}\mathcal{O}_{\mathcal{B}_{\Phi}}\rightarrow\operatorname{Lie}G^{\sharp}\rightarrow\operatorname{Lie}B\rightarrow 0.$